1.In the equation (𝑃+π‘Ž/𝑉2)(π‘‰βˆ’π‘)=π‘π‘œπ‘›π‘ π‘‘π‘Žπ‘›π‘‘, determine the units of a and b.

2. The square of the speed v of an object undergoing acceleration a is a variable of a function a and the displacement S according to the expression given by

𝑣2=π‘˜π‘Žπ‘šπ‘†π‘›

What dimensions should k have in order for the expression to be dimenssionally consistent?


1
Expert's answer
2021-03-10T17:22:12-0500

1)


"[b]=[L^3]=m^3\\ [a]=Pa(m^3)^2=Nm^4"

2)


"𝑣^2=𝑘𝑎^𝑚𝑆^𝑛\\ [𝑣]^2=[𝑎]^𝑚[𝑆]^𝑛\\ [LT^{-1}]^2=[LT^{-2}]^𝑚[L]^𝑛\\ -2=-2m\to m=1\\ 2=m+n=1+n\to n=1"

k will be dimensionless constant.